The second angle of a triangle is 18 degrees more than twice the first. The third angle is three times as large as the second. How large are the angles?
Answer by Nate(3500)

First angle = x
Second angle = 2x + 18
Third angle = 3(2x + 18) = 6x + 54
x + 2x + 18 + 6x + 54 = 180
9x + 72 = 180
9x = 108
x = 12
First angle = 12 degrees
Second angle = 42 degrees
Third angle = 126 degrees